A capacitor consists of two square metal plates, each measuring 5.00×10^2m on a side. In between the plates is a sheet of mica measuring 1.00×10^4m thick.


(a) What is the capacitance of this capacitor? If the charge in one plate is 2.00×10^8C.


(b) What is the potential difference?


(c) What is the electric field between the plates?


Give: Permittivity of Mica (E= 4.8×10^-11C²/N.m²)

1
Expert's answer
2022-03-24T11:01:41-0400

(a) What is the capacitance of this capacitor?


"C=\frac{\epsilon A}{d}\\
=\frac{4.8\times10^{-11}\times 5.00\times 10^{-2}}{1.00\times 10^{-4}}=2.40\times 10^{-8}\:\rm F"

(b) What is the potential difference between plates


"V=\frac{q}{C}\\
=\frac{2.00\times 10^{-8}}{2.40\times 10^{-8}}=0.833\:\rm V"

(c) What is the electric field between the plates?


"E=\frac{V}{d}\\
=\frac{0.833}{1.00\times 10^{-4}}=8330\:\rm V/m"